OpenStudy (anonymous):

What is the common difference of a 57-term arithmetic sequence where the first term is -25 and the sum is 17,727?

OpenStudy (anonymous):

17727/57=311 -25+28n=311 28n=336 n=12

OpenStudy (anonymous):

THank you, can you help me wiht one more?

OpenStudy (anonymous):

ok. sure

OpenStudy (anonymous):

What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is -10?

OpenStudy (anonymous):

a1=123 a2=113 a3=103 ... (123+13)(123-13+10)/20=816

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

welcome :)